OpenStudy (nautika27):

determine whether the ratios are proportional 2/4 20/36

OpenStudy (rcp031):

2/4=1/2 & 20/36=5/9 hence 2/4 is not proportional to 20/36

OpenStudy (mathmale):

One approach would be to reduce both 2/4 and 20/36 to the lowest possible terms. Are the results equal?